Using the formula A = P(1 + r)n  where A is the future value of the investment, P is the principal, r is the fixed annual interest rate, and n is the number of years, how many years will it take an investment to double if the interest rate per annum is 20%?


Solution:




1
Expert's answer
2021-11-29T13:18:20-0500
"A=P(1+r)^n=2P"

"(1+r)^n=2"

"\ln((1+r)^n)=\ln(2)"

"n\ln(1+r)=\ln(2)"

"n=\dfrac{\ln 2}{\ln (1+r)}"

Given "r=0.20"


"n=\dfrac{\ln 2}{\ln (1+0.2)}"

"n\approx3.8\ years"